OpenStudy (anonymous):

The sum of 4y feet and 2i inches equals (A) 6iy feet (B) (12y+i/6) ft (C) 12 1/6 ft (D) (12y+24i) ft (E) 12 1/6 iy ft

OpenStudy (anonymous):

you have to either convert feet to inches or inches to feet

OpenStudy (anonymous):

soo...4y + 1/16i?

OpenStudy (anonymous):

i would have said \(4y+\frac{1}{6}i\) feet but i don't see that answer

OpenStudy (anonymous):

Wait. Meant 1/36i

OpenStudy (anonymous):

1 inch is 1/12 of a foot, so 2 inches is 2/12 = 1/6 of a foot

OpenStudy (anonymous):

But it says yards..

OpenStudy (anonymous):

what says yards?

OpenStudy (anonymous):

the "sum of 4y yards"